OpenStudy (anonymous):

multiply and simplify: 5^(-8)∙5^9

OpenStudy (anonymous):

5^(-8) * 5^(9) Use rules of exponents 5^(-8+9) = 5^1 = 5

OpenStudy (anonymous):

is that the same if the the question is 5 to the -8 power times 5 to the 9th power?

OpenStudy (anonymous):

Yes same answer